OpenStudy (anonymous):

I've searched a lot about these Elliptic curves...but i cant find a suffice answer :/ Idea of this question comes from workin on this problem : What kind of number will result from \(\sqrt{n+\sqrt[3]{n+1}}\) where \( n \in \mathbb{N}\).

OpenStudy (anonymous):

there are some positive integers that makes it integer... \(n=7 \Rightarrow \sqrt{7+\sqrt[3]{7+1}}=2 \) \(n=2196 \Rightarrow \sqrt{2196+\sqrt[3]{2196+1}}=47 \) this got me thinking on how to find all positive integers that makes \(\sqrt{n+\sqrt[3]{n+1}}\) an integer. ------------------------------------------------------------------- let \(m=\sqrt{n+\sqrt[3]{n+1}}\) First of all... in order for \(m\) to be an integer, also \(n+\sqrt[3]{n+1}\) has to be integer, and therefore \(\sqrt[3]{n+1}\) has to be integer. Let's call this number \(a\). We have (because \(a=\sqrt[3]{n+1}\) ) : \(n=a^3-1\) And substituting in the original equation we get: \(m^2=a^3+a-1\) Now solving the original equation is equivalent to solve this Diophantine (considering that \(m\) is positive...) . But there is a problem: the equation found is an elliptic curve...

OpenStudy (anonymous):

this means that we have to find all the points having integer coordinates on this elliptic curve.(o.O) I'm sorry, I can't go further...

OpenStudy (anonymous):

@KingGeorge @Neemo @eliassaab

OpenStudy (anonymous):

what math are you in

OpenStudy (anonymous):

this is something related to diophantine equations

OpenStudy (anonymous):

saura that problem was a result of workin on this

OpenStudy (anonymous):

Oh... ya

OpenStudy (anonymous):

I can't help you with the algebra (I hate number theory more than nearly anything else), but if you just want a lot of pairs to play around with I can very quickly write a program to find them...

OpenStudy (anonymous):

:) thank u very much...actually i want to know if there is an analyticall approach or not.

OpenStudy (anonymous):

I've been reading some of the literature on this...(a) there are only finitely many solutions.

OpenStudy (anonymous):

There is an algorithm by Alan Baker to find all of them.

OpenStudy (anonymous):

so there are finite solutions for that...is that for all elliptic curves like this :\[y^2=x^3+Ax+B\]?

OpenStudy (anonymous):

Yep. Those are called Weierstrass functions (don't confuse this with his everywhere continous but nowhere differentiable function).

OpenStudy (anonymous):

I'm going to assume you know at least a little group theory. I don't know about finding integer solutions, but the rational solutions (as points) form a group, with composition of two points defined as the third point of intersection of the line they form and the original function. You can even find the tangent line at a single point, and where it intersects the graph again will be a rational solution (essentially P+P). I don't quite understand why this works...but it does. In any event, this gives you a (painful) way to get solutions, and you could in theory find the entire group (which you could then prove is the entire group by finding every composition).

OpenStudy (anonymous):

It's also finitely generated...so you could find the generating set (no idea how to, though).

OpenStudy (anonymous):

This is all coming from a paper someone wrote (see attached).

OpenStudy (anonymous):

Thank you so much for all your help I really appreciate it :)

OpenStudy (anonymous):

now i must say i really want to go into group theory

OpenStudy (anonymous):

It's a wonderful subject. I find it much more fun than analysis or number theory. You could do worse than to get Algebra by Michael Artin...it's a very good textbook on Linear Algebra and group theory.

OpenStudy (anonymous):

Nice approach @mukushla

OpenStudy (anonymous):

oh ... thank u...but the main part is still a mystery

OpenStudy (experimentx):

*